10x-25+10x+3x^2+2x^2=360
Answer by checkley77(12844)

10x-25+10x+3x^2+2x^2=360
5x^2+20x-25-360=0
5x^2+20x-385=0
5(x^2+4x-77)=0
5(x+11)(x-7)=0
x+11=0
x=-11 ans.
x-7=0
x=7 ans.